Give first the step you will use to separate the variable and then solve the equation:
z/3 = 5/4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Now we have to multiply both sides of the equation by 3,
Then we get,
= (z/3) × 3 = (5/4) × 3
= x = 15/4

Thanks